A chocolate factory has 2 kinds of chocolate minibar which cost 14 and 8 per kg., respectively. How many kgs. each must be used to make a mixture of 42 kg to sell at 10 per kg.?
Answer by TimothyLamb(4379)

x = kg of $14 bars
y = kg of $8 bars
---
x + y = 42
14x + 8y = 10*42
---
put the system of linear equations into standard form
---
x + y = 42
14x + 8y = 420
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = kg of $14 bars = 14
y = kg of $8 bars = 28
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---